consider the following quadratic equation: 2x^2-40x+150=0. a trusted mathematician tells us that the roots for this equation are 15 and 75. however, when you try to solve it, the roots turns out to be 15 and 5. explain why 75 could be a solution while 5 is not


1
Expert's answer
2021-07-12T17:18:43-0400

The given quadratic equation

"\Rightarrow 2x^2-40x+150=0"

Solving the quadratic equation,

"\Rightarrow x^2-20x+75=0"

"\Rightarrow x^2-15x-5x+75=0"

"\Rightarrow (x-15)(x-5)=0"

So, from the above we can conclude that x = 15 and x= 5 are the roots of the equation.

There is a chance that there is some calculation mistake in the solving the quadratic equation.